Write a quadratic equation for -7,5
Answer by nerdybill(7384)

I'm assuming you want a quadratic that crosses at a point (x,y)
where
(x,y) = (-7,5)
.
If so, consider:
y = x^2 + x + C
Plug in your values, solve for C:
5 = (-7)^2 + (-7) + C
5 = 49 - 7 + C
5 = 42 + C
-37 = C
So, your equation would be:
y = x^2 + x - 37